OpenStudy (anonymous):

Find an equation for the line that is described. Write the answer in two forms. y=mx+b and Ax+By+C=0 Is perpendicular to x-y+2=0 and passes through (3, 1)

Directrix (directrix):

x - y + 2 = 0 x + 2 = y y = x + 2 slope is 1 because 1 is the coefficient of x in this y = mx + b form of a line's equation

Directrix (directrix):

The slope of the perpendicular is -1 which is the negative reciprocal of 1, the slope of the first line.

Directrix (directrix):

Perpendicular at (3,1) y - 1 = (-1) ( x - 3) y - 1 = -x + 3 y = -x + 4 ---> Equation of perpendicular in y = mx + b form y + x - 4 = 0 1x + 1y -4 = 0 ----> Equation of perpendicular in Ax + By + C = 0 form.

OpenStudy (anonymous):

thank you again

Directrix (directrix):

Glad to help.